# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ III.3 DS-$\mathbb{Z}_{4}$ Code
The domain wall between the DS and $\mathbb{Z}_{4}$ condenses $\{1,e^{2}m^{2}\}$ of the $\mathbb{Z}_{4}$ anyons, while all anyons of DS can propagate into the $\mathbb{Z}_{4}$. In contrast to our prior examples, involving a single TQFT, the $\mathcal{L}_{\text{$\mathbb{Z}_{4}$-DS}}$ domain wall is non-invertible. This implies that particles do not tunnel across symmetrically. For example, the $e$ anyon cannot tunnel into the DS phase as it is confined. This is natural since the two phases have a different number of anyons. We defer the details of how this domain wall is constructed, and its relation to anyon condensation, to Ex. IV.5. Since the wall is non-invertible, its action cannot be read from the generators $\{e,m\}$ alone. A more complete description of domain walls and their relations to degeneracy is discussed in Sec. VI. For illustration, we now give the DW’s action on each anyon of $\mathbb{Z}_{4}$.
@@
The confined anyons inside the DS phase require an energy that is linearly proportional to the distance they are translated. As such, they are not elementary excitations of the model anymore. The different behaviors of anyons, when tunneling through or reflecting off a non-invertible wall, will be important for error correction. We observe that new logical operators are worldlines of the anyons which condense, the $e^{2}m^{2}$s, on the $DS$ patches. Let us fix a certain $DS$ patch (upper right) and then connect it to other patches, forming a spanning tree as illustrated in Fig. 8. The $e^{2}m^{2}$ strings can also be thought of as one $e$ loop and one $m$ loop between the two patches.
With four patches Ex. V.9 gives $\operatorname{GSD}=8$. Note that other choices are possible for the strings since any closed loop of strings between the patches is trivial. These strings commute and square to one $e^{4}m^{4}=1$. We therefore identify them with the Pauli $\widetilde{X}_{i}$ operators on each of the three logical qubits. In addition to patches exchanging $e^{2}m^{2}$s, we can also form non-contractible loops, of $e$ or $m$ world-lines, around each patch, which are confined in the DS phase, and they cannot therefore be contracted to trivial loops inside the patch. These loops are depicted in Fig. 8. Interestingly, the loops again square to unity, this time because the $e^{2}$ anyons are deconfined inside $DS$ and the, e.g., $e^{2}$, loop can be contracted. Further, the $e$ or $m$ world-line loops anticommute with the Pauli $\widetilde{X}_{i}$ operators as $e^{2}m^{2}$ and $e$ anticommute Eq. (12). They can then be used to define the Pauli $\widetilde{Z}_{i}$ logical operators acting on each new logical qubit.
@@
Figure 8: Hybrid three-logical-qubit $\mathbb{Z}_{4}$-DS patch code comprising of four DS patches inside a $\mathbb{Z}_{4}$ bulk. The logical $\widetilde{Z}$ operators consist of $e$ loops (solid lines) around three patches. The logical $\widetilde{X}$ operators are an $e^{2}m^{2}$ (22 solid lines and 22 dashed lines) string between three patches and one patch. These patches encode 33 qubits in addition to possible extra $\operatorname{GSD}$ from the boundaries. For lattice-level details of the boundaries, see Fig. 32 and supporting text.
III.3.1 Patch Error Rates
It has been demonstrated that the subthreshold behavior of a surface code depends critically on its precise shape and boundary conditions [55]. We now discuss how, composite algebraic dimension and the resulting logical operators provide another route to tailor the code’s performance.
Let us first compare the probabilities of logical errors of qubits, defined with respect to DS patches, with the $d=4$ qudits defined via $\mathbb{Z}_{4}$ punctures or domain walls. Assume that $X,X^{\dagger},Z,$ and $Z^{\dagger}$ errors each occur with a probability $p$. The left panel of Fig. 4 shows how, for a logical error to occur, an $m$ particle must be exchanged by smooth holes by a string of $X$ or $X^{\dagger}$ operators. Alternatively, an $e,\bar{e}$ particle-anti-particle pair must be created, encircle a puncture via a loop of $Z$ operators, and self-annihilate. Since these particles are each deconfined, each event may be seeded by a single error which occurs with probability $p$. To simplify analysis, we will not concern ourselves with the likelihood that they take a particular path and take it to be unity. Since there is some probability that the particles self-annihilate or take alternate paths, our rough estimate upper bounds the probability of a logical error.
On the other hand, as illustrated in Fig. 8, the error rates of the DS patch logical qubit operators are both higher order and anisotropic. Here, creating a charge or flux particle-particle pair, encircling a patch, and re-annihilating constitutes a logical Pauli-$\widetilde{Z}$ operator. Since it involves a single deconfined $Z_{4}$ anyon generator, it is seeded by an error occurring with probability $p$. On the other hand, the other $\widetilde{X}$ operators are realized by exchanging $e^{2}m^{2}$ particles between a pair of DS patches. Since to generate these particles one must act by $X^{2}$ and $Z^{2}$ simultaneously, the overall probability for such an anyon to be generated by an error is $p^{4}\ll p$ for $p<1$. A string consisting of $X^{2}$ and $Z^{2}$ will further be required to mediate the quasiparticle transport between the patches. Again, our analysis skips over details, including the fact that it is unlikely for all four (two $e$ and two $m$) particles to collectively randomly transfer between patches. Instead, given random errors, the particles are expected to disperse radially in different random directions. This random scattering of individual two $e$ and two $m$ particles, which can be viewed as a brownian motion of the errors, will reduce the error rate as only $e^{2}m^{2}$ condenses at the DS-$\mathbb{Z}_{4}$ domain wall. This anisotropy leads one to conclude that, by condensing DS patches or puncturing holes, one may take advantage of a tradeoff between logical Hilbert space dimension and logical error rate [55] to best suit a particular hardware architecture.
Figure 9: Left: a logical-qubit DS-$\mathbb{Z}_{4}$ surface code on a square lattice with side length NN, the thin rectangular DS slabs have width $s$. Right: logical-qubit DS-$\mathbb{Z}_{4}$ surface code with triangular DS patches with side length $t$.
For concreteness, let’s fix a square with sides of length $N$ supporting a lattice of $2N^{2}$ qudits ($N^{2}$ unit cells each with 2 qudits). One can put two rectangular DS patches at the two ends of the square with a $\mathbb{Z}_{4}$ phase. The patches are separated by smooth (or rough) boundaries on top and bottom as shown in the left panel of Fig. 9. For consistency, we will compute all code properties in terms of qubit resources. This quantum code (for $s<3N/8$) has
$\llbracket\,n,\,k,\,d\rrbracket=\llbracket\,4N^{2},\,1,\,2N\rrbracket$ (32)
where n is the number of physical qubits ($2$ per qudit), $k$ is the number of logical qubits and $d$ is the weight of the smallest logical operator ($\widetilde{Z}$ in this case) [76, 2, 77] where $N$ qudits or $2N$ qubits are involved. Extending the notation, to denote the distance of X (first) and Z (second) logical operators, we write
$\llbracket\,n,\,k,\,d_{X},\,d_{Z}\rrbracket=\llbracket\,4N^{2},\,1,\,8(N-2s),\,2N\rrbracket$ (33)
This expression highlights how this code exhibits a biased error rate [55].
Alternatively, to tune this error rate, the two patches can be placed at the two corners as shown in the right panel of Fig. 9. The parameter $t$ is a free geometric parameter that will affect the code distance:
$\llbracket\,n,\,k,\,d_{X},\,d_{Z}\rrbracket=\llbracket\,4N^{2},\,1,\,\lfloor 8\sqrt{2}(N-t)\rfloor,\,\lfloor 2\sqrt{2}t\rfloor\rrbracket$ (34)
This is evidently valid for the continuum limit. For a square lattice we have instead:
$\llbracket\,n,\,k,\,d_{X},\,d_{Z}\rrbracket=\llbracket\,4N^{2},\,1,\,16(N-t-1),\,4(t+1)\rrbracket$ (35)
[…]
Figure 10: Comparison between $\widetilde{X}$, $\widetilde{Z}$ logical error distances $d_{X}$ and $d_{Z}$ Eq.(35) of the DS-$\mathbb{Z}_{4}$ triangular patches with different dimension $t$ in Fig. 9 and the symmetric error distance of the surface code $d_{\text{S.C.}}$.
Lastly, in addition to new qubits defined via patches, we may also, via punctures, define a set of qubits inside the DS. This generalizes the result of Ref. 40, which only defined the DS toric code. Here, the error correction situation is even better due to confinement. In this case, denote the $\widetilde{X}$ operator as one that exchanges a boson between two punctures with the unique DS boundaries. In analogy, the $\widetilde{Z}$ operator is a (anti-) semion’s world-line loop operator. From the perspective of the $\mathbb{Z}_{4}$ construction, both of these particles are generated by the product of two $\mathbb{Z}_{4}$ Pauli operators.
Specifically, boson-anti-boson pairs are generated by local $X^{2}$ or $Z^{2}$ errors and (anti-) semions are minimally generated by a ($Z^{\dagger}$) $Z$ and $X$ pair of operators. Hence, these errors each come with probability $p^{2}\ll p$ for $p\ll 1$. Further, within the DS phase, the errors that occur with probability $p$ generate confined particles. Since they are confined, i.e. they are not mobile and have an energy cost linear in the particle’s worldline, these errors will be vastly simpler to correct compared to the deconfined errors in the $D(Z_{2})$ codes. Again, this is because the particles entire worldline can be extracted via syndromes instead of just the worldlines’ endpoints. Finally, not that this extra simplification in decoding confined particles holds in the yellow DS regions in Fig. 10. As a result, while the Hilbert space overhead is larger, such codes are worthwhile to pursue because they dramatically simplify the decoding and recovery stages of quantum error correction, which are currently the bottleneck to topological codes’ scalability.
IV Microscopic Boundary Formulation
@@
The classification of the boundaries and domain walls of a twisted quantum double based on the (not necessarily Abelian) group $G$ is given by the Lagrangian subgroups of the topological order [46, 48]. As seen in Sec. III.1.2, the set of mutually compatible condensing anyons completely specifies a boundary. In the case of untwisted quantum doubles, these Lagrangian subgroups are in one-to-one correspondence with a subgroup $K\subseteq G$ and a function $\omega:K\times K\to U(1)$ that associates a $U(1)$ phase to each ordered pair of elements in the subgroup $K$. This function is a 2-cocycle since it obeys:
@@
In the case of twisted quantum doubles extra constraints hold for the boundary [49]. For example, the DS phase is based on the group $\mathbb{Z}_{2}$, which has two subgroups, namely the trivial subgroup and the whole group. However, since it is a twisted quantum double with a non-trivial 3-cocycle, it cannot have a boundary with a non-trivial subgroup [49].
@@
In what follows, we will not deal with the classification in terms of cocycles directly. Instead, we will only use the physically intuitive Lagrangian subgroup of anyons explicitly while the cocycle picture will be implicit. This is contrary to Refs. 45, 49, 47 that abstractly formulated how to construct microscopic lattice models for a (twisted) quantum double’s gapped boundary or domain walls using cocycles. In the case of quantum doubles, they are formed by defining subgroup variants of the star and plaquette operators (5). The modified star operators have a modified action that associates a $U(1)$ phase depending on the qudits’ values. Consequently, they are not readily implementable experimentally. This issue will arise for $1D$ domain walls and $0D$ defects that will need an extra $U(1)$ phase (those with a nontrivial 2-cocycle) [45, 49, 47]. From another perspective, these abstract terms while correct, are not derived from the bulk terms.
Other constructions were proposed recently to obtain Pauli stabilizers for the boundaries of Abelian (twisted) quantum doubles [50]. One important difference is that our algorithm also produces the 0D defects. In addition, it produces topologically complete models, meaning that the only local operators that commute with the stabilizers are products of other stabilizers [54, 50]. Other choices for the initial (twisted) quantum double stabilizers are still valid for our algorithm. For example, spatially disconnected stabilizers introduced in Ref. 50, will not require extra steps to ensure topological completion in our construction.
IV.1 Boundaries by Local Condensation
The operators presented in the previous section, while sufficient to mathematically study the properties of each boundary, do not provide a simple prescription for experimental implementation. To this point, for Abelian twisted quantum doubles, a description of the bulk in terms of generalized Pauli stabilizers was recently provided [40]. However, a simple construction of Pauli stabilizers for the boundaries and domain walls was still lacking. Condensation of anyons is the suitable tool to study boundaries [46, 78, 79]. Inspired by the correspondence between the classification of boundaries and Lagrangian subgroups of the bulk, we present an algorithm to construct different boundaries, domain walls, and $0D$ defects of Abelian topological orders described by (twisted) quantum doubles using local condensation. Our construction of the boundaries is similar to the procedure for condensing the bulk of untwisted quantum double to produce twisted quantum doubles [40], but with two modifications. First, we condense a Lagrangian group instead of a partial condensation that leaves non-trivial topological order afterward. Second, we only condense a region in space that leaves the boundary terms we are interested in.
@@ IV.1 Boundaries by Local Condensation
Local condensation–Condensing a set of anyons globally, whether a Lagrangian group or one of its subgroups, will switch the topological phase into another one. However, if the condensation is carried over locally, we end up with two phases with a domain wall at their interface. If one of the phases is vacuum, then the interface will be a boundary (maximal condensation); otherwise, it will be a domain wall between two phases (partial condensation). In the $\mathbb{Z}_{4}$ example, condensing $\mathcal{L}_{rough}=\{1,e,e^{2},e^{3}\}$ over a region $R$ of space will result in vacuum in that region. Additionally, the boundary $\partial R$ will correspond to the rough boundary of $\mathbb{Z}_{4}$ Eq. (13). Similarly, condensing $e^{2}$ over a region $R$ will result in $\mathbb{Z}_{2}$ over that region. The boundary $\partial R$ will be a domain wall between $\mathbb{Z}_{4}$ and $\mathbb{Z}_{2}$. Condensing $m^{2}$ over $R$ would have resulted also in the $\mathbb{Z}_{2}$ phase over that region. However, it will correspond to a different domain wall between $\mathbb{Z}_{4}$ and $\mathbb{Z}_{2}$. The correspondence between anyon condensation and boundaries ensures that all domain walls can be constructed this way.
The takeaway from the $\mathbb{Z}_{4}$ example is that condensing anyons over a region $R$ should be enough to construct all possible boundaries at $\partial R$. Since any domain wall between two phases is a boundary of a related phase, we will focus on boundaries without loss of generality. We will treat concrete examples of domain walls in Sec.IV.4 and finite domain walls ending with 0D defects in App. B.
IV.2 Algorithm to Construct Boundary Stabilizers

### v4

@@ IV.4.4 Hybrid DS-$\mathbb{Z}_{4}$ code results
We still expect that the logical error rate will be lower than that of the planar code, as the average weight of the logical $\widetilde{X}$ operators is now higher. The domain wall between the DS and $\mathbb{Z}_{4}$ condenses $\{1,e^{2}m^{2}\}$ of the $\mathbb{Z}_{4}$ anyons, while all anyons of DS can propagate into the $\mathbb{Z}_{4}$. In contrast to our prior examples, involving a single TQFT, the $\mathcal{L}_{\text{$\mathbb{Z}_{4}$-DS}}$ domain wall is non-invertible. This implies that particles do not tunnel across symmetrically. For example, the $e$ anyon cannot tunnel into the DS phase as it is confined. This is natural since the two phases have a different number of anyons. We defer the details of how this domain wall is constructed, and its relation to anyon condensation, to Ex. V.5. Since the wall is non-invertible, its action cannot be read from the generators $\{e,m\}$ alone. A more complete description of domain walls and their relations to degeneracy is discussed in Sec. VI.2. For illustration, we now give the DW’s action on each anyon of $\mathbb{Z}_{4}$.
@@
Figure 12: Layout for a hybrid DS-$\mathbb{Z}_{4}$ code on a plane which has $\operatorname{GSD}=2$. The logical operators $\widetilde{Z}$ are strings of Pauli $Z^{2}$ creating the boson $b$. While the logical $\widetilde{X}$ are strings of Pauli $X$. Alternatively, a string of simultaneous $X^{2}$ and $Z^{2}$ can also create a logical $\widetilde{X}$ between the two DS patches.
One can naturally explore other layouts for these codes. For example, we can push the patches to the bulk of the $\mathbb{Z}_{4}$ as in Fig. 13, also see Ex. VI.7 gives $\operatorname{GSD}=8$.
@@
Figure 13: Hybrid three-logical-qubit $\mathbb{Z}_{4}$-DS patch code comprising of four DS patches inside a $\mathbb{Z}_{4}$ bulk. The logical $\widetilde{Z}$ operators consist of $e$ loops (solid lines) around three patches. The logical $\widetilde{X}$ operators are an $e^{2}m^{2}$ (22 solid lines and 22 dashed lines) string between three patches and one patch.
The patches can exchange $e^{2}m^{2}$ anyons between them. These strings commute and square to one $e^{4}m^{4}=1$. We therefore identify them with the Pauli $\widetilde{X}_{i}$ operators on each of the three logical qubits. In addition to patches exchanging $e^{2}m^{2}$s, we can also form non-contractible loops, of $e$ or $m$ world-lines, around each patch, which are confined in the DS phase, and they cannot therefore be contracted to trivial loops inside the patch. These loops are depicted in Fig. 13. Interestingly, the loops again square to unity, this time because the $e^{2}$ anyons are deconfined inside $DS$ and the, e.g., $e^{2}$, loop can be contracted. Further, the $e$ or $m$ world-line loops anticommute with the Pauli $\widetilde{X}_{i}$ operators as $e^{2}m^{2}$ and $e$ anticommute Eq. (12). They can then be used to define the Pauli $\widetilde{Z}_{i}$ logical operators acting on each new logical qubit.
To summarize this section, we present the thresholds of different codes subject to different noise models and using ILP, BP-OSD-CS, and MWPM-4 decoders in Table 1.
# Codes Noise ILP BP-CS MWPM-4
1 Planar $\mathbb{Z}_{2}$ Pure X 11% 10.75% 10.3%
Depol. 18% 16% 15.25%
2 Planar $\mathbb{Z}_{4}$ Pure X 18% 18% 14%
Depol. 27% 26.5% 17.25%
3 Even Pure X 17.5% 18% 15.5%
Depol. 27% 26% 17.75%
4 DS Pure X 17.5% 21% N/A
Depol. 20.5% 22% N/A
5 DS - $\mathbb{Z}_{4}$ Pure X 18.5% 19% N/A
Depol. 26.5% 26.5% N/A
Table 1: Code capacity thresholds across different noise models and decoders.
The computing resources used to produce the table results are a single node with 64 cores (multicore processing on a single node), with a maximum job run time of 24 hours for the ILP decoder. However, for the latest belief propagation-based decoder, the run times are largely less than 10 hours. We used shots between $5\times 10^{5}$ and $10^{8}$. These results show that using the same number of qudits, one can still achieve different thresholds or logical error rates by virtue of their connectivity. Indeed, as shown in Table 2, the encoding rates and number of qubits (assuming each qudit is made out of two qubits) are comparable for these codes, while they still exhibit different behaviors as evidenced in Table 1. As shown in Fig. 14, the logical error rates are successively suppressed as we change the codes. For instance, both the even and the hybrid code use the same asymptotic number of qudits, however, the hybrid code has a lower logical error rate. This shows that spatial structure, composite-dimensionality and stabilizer connectivity are three valuable resources for topological codes.
Figure 14: Comparison of logical error rates among different codes subject to pure $X$ noise using the belief-propagation algorithm discussed in Sec. IV.3. Solid lines represent $d=3$ while dashed lines represent $d=7$. The plot shows monotonic error suppression between the different codes.
# Codes $n$ $k$ $d$
1 Planar $\mathbb{Z}_{2}$ $2d^{2}$ 1 $d$
2 Planar $\mathbb{Z}_{4}$ $4d^{2}$ 2 $d$
3 Even $4d^{2}$ 2 $d$
4 DS $4d^{2}$ 4 $d$
5 DS - $\mathbb{Z}_{4}$ $4d^{2}$ 2 $d$
Table 2: Code parameters $n$, $k$, and $d$ for each evaluated code family.
V Microscopic Boundary Formulation
To make the results and codes above rigorous, this section derives the lattice models presented and provides an algorithm to do so. Our algorithm takes as input a set of stabilizers and returns a new set of stabilizers describing the new Abelian twisted quantum double, which may contain boundaries and domain walls in addition to bulk terms. We present examples highlighting its use in constructing DS patches embedded in a $\mathbb{Z}_{4}$ parent phase at both a concrete and conceptual level.
Conceptually, we show how the understanding of the classification of boundaries can aid their construction from the bulk Hamiltonian terms. This is manifested in a concrete lattice-level algorithm. The concrete construction of the new codes in the previous section is a by-product of this algorithm. Its utility can extend well beyond that. The intuition and perspective presented here can be especially useful for the recently proposed schemes that use multiple phases to perform universal quantum computing [94, 135].
This section is organized into five subsections. Sec. V.1 lays out the classification of the boundaries and discusses differences from other works, Sec. V.2 explains at a conceptual level why boundaries can be constructed from bulk terms, Sec. V.3 gives the concrete lattice-level algorithm for their constructions, Sections V.4 and V.5 give examples of boundaries and domain walls construction. The construction of $0D$-defects is deferred to App. B.
V.1 Boundaries Classification
@@
The classification of the boundaries and domain walls of a twisted quantum double based on the (not necessarily Abelian) group $G$ is given by the Lagrangian algebras of the topological order [114, 112]. As seen in Sec. III.1.2, the set of mutually compatible condensing anyons completely specifies a boundary. In the case of untwisted quantum doubles, these Lagrangian subgroups are in one-to-one correspondence with a subgroup $K\subseteq G$ and a function $\omega:K\times K\to U(1)$ that associates a $U(1)$ phase to each ordered pair of elements in the subgroup $K$. This function is a 2-cocycle since it obeys:
@@
In the case of twisted quantum doubles, extra constraints hold for the boundary [37]. For example, the DS phase is based on the group $\mathbb{Z}_{2}$, which has two subgroups, namely the trivial subgroup and the whole group. However, since it is a twisted quantum double with a non-trivial 3-cocycle, it cannot have a boundary with a non-trivial subgroup [37].
@@
In what follows, we will not deal with the classification in terms of cocycles directly. Instead, we will only use the physically intuitive Lagrangian subgroup of anyons explicitly, while the cocycle picture will be implicit. This is contrary to Refs. 12, 37, 52, 53 that abstractly formulated how to construct microscopic lattice models for a (twisted) quantum double’s gapped boundary or domain walls using cocycles. In the case of quantum doubles, they are formed by defining subgroup variants of the star and plaquette operators (6). The modified star operators have a modified action that associates a $U(1)$ phase depending on the qudits’ values. Consequently, they are not readily implementable experimentally. This issue will arise for $1D$ domain walls and $0D$ defects that will need an extra $U(1)$ phase (those with a nontrivial 2-cocycle) [12, 37, 52, 53]. From another perspective, these abstract terms, while correct, are not derived from the bulk terms.
Other constructions were proposed recently to obtain Pauli stabilizers for the boundaries of Abelian (twisted) quantum doubles [130]. One important difference is that our algorithm also produces the 0D defects. In addition, it produces topologically complete models, meaning that the only local operators that commute with the stabilizers are products of other stabilizers [29, 130]. Other choices for the initial (twisted) quantum double stabilizers are still valid for our algorithm. For example, spatially disconnected stabilizers introduced in Ref. 130, will not require extra steps to ensure topological completion in our construction.
V.2 Boundaries by Local Condensation
The operators presented in the previous works [12, 37, 52, 53], while sufficient to mathematically study the properties of each boundary, do not provide a simple prescription for experimental implementation. To this point, for Abelian twisted quantum doubles, a description of the bulk in terms of generalized Pauli stabilizers was recently provided [66]. However, a simple construction of Pauli stabilizers for the boundaries and domain walls was still lacking. Condensation of anyons is a suitable tool to study boundaries [114, 38, 105]. Inspired by the correspondence between the classification of boundaries and Lagrangian subgroups of the bulk, we present an algorithm to construct different boundaries, domain walls, and $0D$ defects of Abelian topological orders described by (twisted) quantum doubles using local condensation.
Our construction of the boundaries is similar to the procedure for condensing the bulk of the untwisted quantum double to produce twisted quantum doubles [66], but with two modifications. First, we condense a Lagrangian group instead of a partial condensation that leaves non-trivial topological order afterward. Second, we only condense a region in space that leaves the boundary terms we are interested in.
@@ V.2 Boundaries by Local Condensation
Local condensation–Condensing a set of anyons globally, whether a Lagrangian group or one of its subgroups, will switch the topological phase into another one. However, if the condensation is carried over locally, we end up with two phases with a domain wall at their interface. If one of the phases is vacuum, then the interface will be a boundary (maximal condensation); otherwise, it will be a domain wall between two phases (partial condensation). In the $\mathbb{Z}_{4}$ example, condensing $\mathcal{L}_{rough}=\{1,e,e^{2},e^{3}\}$ over a region $R$ of space will result in vacuum in that region. Additionally, the boundary $\partial R$ will correspond to the rough boundary of $\mathbb{Z}_{4}$ Eq. (13). Similarly, condensing $e^{2}$ over a region $R$ will result in $\mathbb{Z}_{2}$ over that region. The boundary $\partial R$ will be a domain wall between $\mathbb{Z}_{4}$ and $\mathbb{Z}_{2}$. Condensing $m^{2}$ over $R$ would have resulted in the $\mathbb{Z}_{2}$ phase over that region. However, it will correspond to a different domain wall between $\mathbb{Z}_{4}$ and $\mathbb{Z}_{2}$. The correspondence between anyon condensation and boundaries ensures that all domain walls can be constructed this way.
The takeaway from the $\mathbb{Z}_{4}$ example is that condensing anyons over a region $R$ should be enough to construct all possible boundaries at $\partial R$. Since any domain wall between two phases is a boundary of a related phase, we will focus on boundaries without loss of generality. We will treat concrete examples of domain walls in Sec.V.5 and finite domain walls ending with 0D defects in App. B.
V.3 Algorithm to Construct Boundary Stabilizers